# The special unitary groups $SU(2n)$ as framed manifolds

### v1

@@ 2. Tensor product decomposition of $E$
$(r,zv)_{R}=p(R(rz,v))=R(r,zv).$
Here in particular $(r,zv)_{R}$ represents $(0,1)_{R}$ if $r=0$ and $(1,0)_{R}$ if $r=1$. The latter is assumed to be specified as the base point of $S^{2}$ for the reason that it is the image of the identity matrix of $SU(2)$ by $p$.
@@
(6) $\begin{split}R^{\{i\}}_{j}(r_{i;j}zz^{\prime},v_{i;j})d(z)&=R^{\{i\}}_{j}(r_{i;j}z^{\prime}_{,}z^{i}v_{i;j})\quad(j\geq 0),\\ R_{j}(r_{i;j}z_{i}z^{\prime}_{i},v_{i;j})d(z_{1})\cdots d(z_{m_{j}})&=\begin{cases}R_{j}(r_{i;j}z^{\prime}_{s},z_{i}^{i}v_{i;j})&(j\geq 1),\\ R_{j}(r_{i;j}z^{\prime}_{i},z_{i-1}^{n}z_{i}^{i}v_{i;j})&(j=0)\quad(\text{where}\ z_{-1}=1).\end{cases}\end{split}$
From (4) we see that $p(R^{\{i\}}_{j}(r_{i;j}z,v_{i;j}))=p(R_{i;j}(r_{i;j},z^{i}v_{i;j}))$ holds for $p:G\to G/S$. This permits us to embed $S^{2}_{i;j}$ in $G/S$ by means of the injective map $(r_{i;j},w_{i;j})_{R}\to p(R^{\{i\}}_{j}(r_{i;j}z,v_{i;j}))$ where $S^{2}_{i;j}=S^{2}$ denotes the 2-sphere consisting of $(r_{i;j},w_{i;j})_{R}$. From (6) we also see that
@@
Let us write $(S_{j}^{2})^{m_{j}}=S^{2}_{1;j}\times\cdots\times S^{2}_{m_{j};j}$ and let $\phi_{j}:(S_{j}^{2})^{m_{j}}\to G/S$ be the map given by $x_{j}=(x_{1;j},\cdots,x_{m_{j};j})\to p(R_{j}(r_{i;j},w_{i;j}))$ where $x_{i;j}=(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j}$. Let us also write $(S^{2})^{(m^{2}-m)/2}=(S_{0}^{2})^{m_{0}}\times\cdots\times(S_{m-2}^{2})^{m_{m-2}}$ and let $\phi:(S^{2})^{(m^{2}-m)/2}\to G/S$ be the map defined by $\phi(x)=\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})$ where $x=(x_{0},\cdots,x_{m-2})$ with $x_{j}\in(S_{j}^{2})^{m_{j}}$. Let us put $R(r_{i;j}z_{i},v_{i;j})=\prod_{j=0}^{m-2}R_{j}(r_{i;j}z_{i},v_{i;j})$. Then, as in the case of $P^{\{i\}}_{j}$, using (5), (5b), (6),
@@
forms the total space of a principal $S$-bundle endowed with the projection map $q:P\to(S^{2})^{(m^{2}-m)/2}$ such that $\phi\circ q=p|P$. Let us write $L_{j}^{m_{j}}=L_{1;j}\boxtimes\cdots\boxtimes L_{m_{j};j}$ and also put $L^{(m^{2}-m)/2}=L_{0}^{m_{0}}\boxtimes\cdots\boxtimes L_{m-2}^{m_{m-2}}$. Then $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong(P,q,(S^{2})^{(m^{2}-m)/2})$ for the reason similar to (7) and therfore we have
@@
where the right hand side denotes the induced bundle of $(G,p,G/S)$ by $\phi$.
@@
By assumption $\phi(x)=\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})$ and $\phi_{j}(x)=p(R_{j}(r_{i;j},w_{i;j}))$. Here $R_{j}(r_{i;j},w_{i;j})$ can be written in the form $R_{j}(r_{i;j},w_{i;j})=\mathrm{diag}(I_{j},M_{j})$ where the first column of $M_{j}$ consists of $\underline{a}_{0;j}$, $\underline{a}_{1;j}$, $\ldots$, $\underline{a}_{m-j-1;j}$ given in (3).
Suppose $\phi(x)=\phi(x^{\prime})$, that is, $\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})=\phi_{0}(x^{\prime}_{0})\cdots\phi_{m-2}(x^{\prime}_{m-2})$. Then due to the definition this can be interpreted as meaning that
$R_{0}(r_{i;0},w_{i;0j})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{0}(r^{\prime}_{i;0},w^{\prime}_{i;0j})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
where we denote by attaching to an element accompanied by $x$ its corresponding element accompanied by $x^{\prime}$. In view of the form of $R_{j}(r_{i;j},w_{i;j})$ we therefore have
@@
Substituting the result of (3) into these equations we have $r_{1;0}\cdots r_{m-1;0}=r^{\prime}_{1;0}\cdots r^{\prime}_{m-1;0}$, $r_{s+1;0}\cdots r_{m-1;0}\bar{w}_{s;0}=r^{\prime}_{s+1;0}\cdots r^{\prime}_{m-1;0}\bar{w}^{\prime}_{s;0}$ $(1\leq s\leq m-1)$ and $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ where $r_{s;0}^{2}+|w_{s;0}|^{2}=1$ and $r_{s;0}\geq 0$. For convenience we assume that $r_{i;0}>0$ and moreover proceed by induction on $i$ in reverse order. First from $\underline{a}_{m-1;0}=\underline{a}^{\prime}_{m-1;0}$ we have $w_{m-1;0}=w^{\prime}_{m-1;0}$ and so it follows that $r_{m-1;0}=r^{\prime}_{m-1;0}$. Next from the second equation $\underline{a}_{m-2;0}=\underline{a}^{\prime}_{m-2;0}$ we see that $r_{m-1;0}w_{m-2;0}=r^{\prime}_{m-1;0}w^{\prime}_{m-2;0}$. But since $r_{m-1;0}=r^{\prime}_{m-1;0}>0$ we have $w_{m-2;0}=w^{\prime}_{m-2;0}$ and so similarly it follows that $r_{m-2;0}=r^{\prime}_{m-2;0}$. Repeating this procedure we obtain subsequently $w_{m-3;0}=w^{\prime}_{m-3;0}$, $r_{m-3;0}=r^{\prime}_{m-3;0}$, $w_{m-4;0}=w^{\prime}_{m-4;0}$, $r_{m-4;0}=r^{\prime}_{m-4;0}$$\cdots$, $r_{2;0}=r^{\prime}_{2;0}$ and finally we obtain $w_{1;0}=w^{\prime}_{1;0}$, $r_{1;0}=r^{\prime}_{1;0}$. At this point we conclude that $x_{0}=x^{\prime}_{0}$ and so the assumption above can be rewritten as
$\phi_{1}(x_{1})\cdots\phi_{m-2}(x_{m-2})=\phi_{1}(x^{\prime}_{1})\cdots\phi_{m-2}(x^{\prime}_{m-2}).$
From this by looking at the form of $R_{j}(r_{i;j},w_{i;j})$ again we have $\underline{a}_{0;1}=\underline{a}^{\prime}_{0;1},\ldots,\underline{a}_{m-2;1}=\underline{a}^{\prime}_{m-2;1}$. Replacing the previous equations by these ones and the condition $r_{i;0}>0$ by $r_{i;1}>0$ in the argument there we can get $x_{1}=x^{\prime}_{1}$, hence the rewritten assumption above is also rewritten as
$\phi_{2}(x_{2})\cdots\phi_{m-2}(x_{m-2})=\phi_{2}(x^{\prime}_{2})\cdots\phi_{m-2}(x^{\prime}_{m-2}).$
Repeating this step we obtain $x_{2}=x^{\prime}_{2}$, $\ldots$, $x_{m-2}=x^{\prime}_{m-2}$ via inductive reduction, which implies $x=x^{\prime}$. This proves the lemma.
@@ 3. Proofs of Theorem and Proposition
Let $\mu:T_{k}\to S^{2}$ be the map defined by $\mu(e^{\eta i},ze^{\theta i})=(-\cos\mu,ze^{\theta i}\sin\mu)$ where $z_{2k-1}=e^{\eta i}$ $(0\leq\eta\leq\pi)$ and $z_{2k}=e^{\theta i}$ $(0\leq\theta<2\pi)$. Taking into account the fact that a principal circle bundle over $S^{1}$ is trivial we see that the classifying map of $p_{k}$ factors through $S^{2}$ where the restriction of $p_{k}$ to $\{1\}\times S^{1}\subset T_{k}$ is viewed as being trivial. Then we have
@@
$(Q_{k},p_{k},T_{k})\cong\mu^{*}(SU(2),p,S^{2})$  $(1\leq k\leq m/2-1)$.
@@
In order to prove this it suffices to show that there is a bundle map covering $\mu$. From the fact that $\mu$ can be rewitten as $\mu(e^{\eta i},ze^{\theta i})=(\cos(\eta/2),ze^{\theta i}\sin(\eta/2))_{R}$ we see that the map $\tilde{\mu}:Q_{k}\to SU(2)$ given by $D(e^{\eta i},ze^{\theta i})d(\bar{z})\to(\bar{z}\cos(\eta/2),e^{\theta i}\sin(\eta/2))$ in fact provides the required bundle map. ∎
@@
$\psi(x,y)=\phi_{0}(x_{0})\cdots\phi_{m-2}(x_{m-2})\psi_{1}(y_{1})\cdots\psi_{n-1}(y_{n-1})$
@@
forms the total space of a principal $S$-bundle endowed with the projection map $\tilde{q}:\tilde{P}\to(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ such that $\psi\circ\tilde{q}=p|P$. Let $\tilde{L}_{k}$ be the complex line bundle over $T_{k}$ associated to $p_{k}$. Write $\tilde{L}^{n-1}=\tilde{L}_{1}\boxtimes\cdots\boxtimes\tilde{L}_{n-1}$ and put $E=L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$. Then we have $(S(E),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong(\tilde{P},q,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})$ and therefore

### v3

@@ 2. Tensor product decomposition of $E$
$(r,zv)_{R}=p(R(rz,v))=R(r,zv)$
where $zv$ can be converted to 1 when $r=0$ because $R(0,zv)\,d(\overline{zv})=R(0,1)$. Thus $(r,zv)_{R}$ represents $(0,1)_{R}$ if $r=0$ and obviously $(1,0)_{R}$ if $r=1$. The latter is assumed to be specified as the base point of $S^{2}$ for the reason that it is the image of the identity matrix of $SU(2)$ by $p$.
@@
(6) $\begin{split}R^{\{i\}}_{j}(r_{i;j}zz^{\prime},v_{i;j})d(z)&=R^{\{i\}}_{j}(r_{i;j}z^{\prime}_{,}z^{i}v_{i;j})\quad(j\geq 0),\\ R_{j}(r_{i;j}z_{i}z^{\prime}_{i},v_{i;j})d(z_{1})\cdots d(z_{m_{j}})&=\begin{cases}R_{j}(r_{i;j}z^{\prime}_{i},z_{i}^{i}v_{i;j})&(j\geq 1),\\ R_{j}(r_{i;j}z^{\prime}_{i},z_{i-1}^{m}z_{i}^{i}v_{i;j})&(j=0)\quad(\text{where}\ z_{0}=1).\end{cases}\end{split}$
From (4) we see that $p(R^{\{i\}}_{j}(r_{i;j}z,v_{i;j}))=p(R_{i;j}(r_{i;j},z^{i}v_{i;j}))$ holds for $p:G\to G/S$. Letting $S^{2}_{i;j}$ denote the 2-sphere consisting of all $(r_{i;j},w_{i;j})_{R}$, this permits us to embed $S^{2}_{i;j}$ in $G/S$ by means of the injective map $(r_{i;j},w_{i;j})_{R}\to p(R^{\{i\}}_{j}(r_{i;j}z,v_{i;j}))$. From (6) we also see that
@@
Let $(S_{j}^{2})^{m_{j}}=S^{2}_{1;j}\times\cdots\times S^{2}_{m_{j};j}$ and $(S^{2})^{(m^{2}-m)/2}=(S_{0}^{2})^{m_{0}}\times\cdots\times(S_{m-2}^{2})^{m_{m-2}}$. Then we have a map $\phi:(S^{2})^{(m^{2}-m)/2}\to G/S$ given by
$x=(x_{0},\cdots,x_{m-2})\to p\biggl{(}\textstyle\prod_{j=0}^{m-2}R_{j}(r_{i;j},w_{i;j})\biggr{)}$
where $x_{j}=(x_{1;j},\cdots,x_{m_{j};j})\in(S_{j}^{2})^{m_{j}}$ with $x_{i;j}=(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j}$. If we put $R(r_{i;j}z_{i},v_{i;j})=\prod_{j=0}^{m-2}R_{j}(r_{i;j}z_{i},v_{i;j})$, then as in the case of $P^{\{i\}}_{j}$, using (5), (5b), (6) we see that
@@
forms the total space of a principal $S$-bundle endowed with the projection map $q:P\to(S^{2})^{(m^{2}-m)/2}$ such that $\phi\circ q=p|P$. If we write $L^{(m^{2}-m)/2}=L_{0}^{m_{0}}\boxtimes\cdots\boxtimes L_{m-2}^{m_{m-2}}$ where $L_{j}^{m_{j}}=L_{1;j}\boxtimes\cdots\boxtimes L_{m_{j};j}$, then $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong(P,q,(S^{2})^{(m^{2}-m)/2})$ for the reason similar to (7). Therfore we have
@@
where $\phi^{*}(-)$ denotes the induced bundle.
@@
By definition $\phi(x)=p(\prod_{j=0}^{m-2}R_{j}(r_{i;j},w_{i;j}))$ and $R_{j}(r_{i;j},w_{i;j})=\textstyle\prod_{i=1}^{m_{j}}R_{i;j}(r_{i;j},w_{i;j})$. Here $R_{j}(r_{i;j},w_{i;j})$ can be written in the form $R_{j}(r_{i;j},w_{i;j})=\mathrm{diag}(I_{j},M_{j})$ where the first column of $M_{j}$ consists of $\underline{a}_{0;j}$, $\ldots$, $\underline{a}_{m-j-1;j}$ given in (3).
Suppose $\phi(x)=\phi(x^{\prime})$, namely $p(\prod_{j=0}^{m-2}R_{j}(r_{i;j},w_{i;j}))=p(\prod_{j=0}^{m-2}R_{j}(r^{\prime}_{i;j},w^{\prime}_{i;j}))$ where we denote by attaching to an element accompanied by $x$ its corresponding element accompanied by $x^{\prime}$. Due to the definition this can be interpreted as meaning that
$R_{0}(r_{i;0},w_{i;0})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{0}(r^{\prime}_{i;0},w^{\prime}_{i;0})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
with $w_{i;j}$ and $w^{\prime}_{i;j}$ converted to 1 when $r_{i;j}=0$ and $r^{\prime}_{i;j}=0$. This attached condition can be verified as in the case of $SU(2)$ as follows. Consider the latter equation of (6) with $z^{\prime}_{i}=1$ and $v_{i;j}$ replaced by $w_{i;j}$ for all $i$. Let $r_{i;j}=0$ for $i=i_{1},\ldots,i_{s}$. Then $|w_{i;j}|=1$ for $i=i_{1},\ldots,i_{s}$, so taking $z_{i}=\bar{w}_{i;j}^{1/i}$ for $i=i_{1},\ldots,i_{s}$ and $z_{i}=1$ except for these $z_{i}$, we have $R_{j}(r_{i;j}z_{i},w_{i;j})d(z_{1})\cdots d(z_{m_{j}})=R_{j}(r_{i;j},z_{i}^{i}w_{i;j})$ where $R_{i;j}(r_{i;j},w_{i;j})=R_{i;j}(0.1)$ for $i=i_{1},\ldots,i_{s}$. But in the case $j=0$, if $r_{i-1;0}\neq 0$, then the above $z_{i}$ needs to be replaced by $z_{i}=w_{i-1;0}^{m}\bar{w}_{i;j}^{1/i}$. Similarly for $w^{\prime}_{s;j}$.
In view of the form of $R_{j}(r_{i;j},w_{i;j})$ we have
@@
Substituting the result of (3) into these equations we have $r_{1;0}\cdots r_{m-1;0}=r^{\prime}_{1;0}\cdots r^{\prime}_{m-1;0}$, $r_{s+1;0}\cdots r_{m-1;0}\bar{w}_{s;0}=r^{\prime}_{s+1;0}\cdots r^{\prime}_{m-1;0}\bar{w}^{\prime}_{s;0}$ $(1\leq s\leq m-2)$ and $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ where $r_{s;0}^{2}+|w_{s;0}|^{2}=1$ and $r_{s;0}\geq 0$. For convenience we assume that $r_{i;0}>0$ and moreover proceed by induction on $i$ in reverse order. First from $\underline{a}_{m-1;0}=\underline{a}^{\prime}_{m-1;0}$ we have $w_{m-1;0}=w^{\prime}_{m-1;0}$ and so it follows that $r_{m-1;0}=r^{\prime}_{m-1;0}$. Next from the second equation $\underline{a}_{m-2;0}=\underline{a}^{\prime}_{m-2;0}$ we see that $r_{m-1;0}w_{m-2;0}=r^{\prime}_{m-1;0}w^{\prime}_{m-2;0}$. But since $r_{m-1;0}=r^{\prime}_{m-1;0}>0$ we have $w_{m-2;0}=w^{\prime}_{m-2;0}$ and so similarly it follows that $r_{m-2;0}=r^{\prime}_{m-2;0}$. Repeating this procedure we obtain subsequently $w_{m-3;0}=w^{\prime}_{m-3;0}$, $r_{m-3;0}=r^{\prime}_{m-3;0}$, $w_{m-4;0}=w^{\prime}_{m-4;0}$, $r_{m-4;0}=r^{\prime}_{m-4;0}$$\cdots$, $r_{2;0}=r^{\prime}_{2;0}$ and finally we obtain $w_{1;0}=w^{\prime}_{1;0}$, $r_{1;0}=r^{\prime}_{1;0}$. At this point we conclude that $x_{0}=x^{\prime}_{0}$ and so the assumption above can be rewritten as
$R_{1}(r_{i;0},w_{i;0})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{1}(r^{\prime}_{i;0},w^{\prime}_{i;0})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
with $w_{i;j}=1$ when $r_{i;j}=0$ and $w^{\prime}_{i;j}=1$ when $r^{\prime}_{i;j}=0$. From this by looking at the form of $R_{j}(r_{i;j},w_{i;j})$ again we have $\underline{a}_{0;1}=\underline{a}^{\prime}_{0;1},\ldots,\underline{a}_{m-2;1}=\underline{a}^{\prime}_{m-2;1}$. Replacing the previous equations by these ones and the condition $r_{i;0}>0$ by $r_{i;1}>0$ in the argument there we can get $x_{1}=x^{\prime}_{1}$, hence the rewritten assumption above is also rewritten as
$R_{2}(r_{i;0},w_{i;0})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{2}(r^{\prime}_{i;0},w^{\prime}_{i;0})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
with the same condition as above. Repeating this step we obtain $x_{2}=x^{\prime}_{2}$, $\ldots$, $x_{m-2}=x^{\prime}_{m-2}$ via inductive reduction, which implies $x=x^{\prime}$. This proves the lemma.
@@ 3. Proofs of Theorem and Proposition
Put $z_{2k-1}=e^{\eta i}$ and $z_{2k}=e^{\theta i}$ for $0\leq\eta,\,\theta<2\pi$ and define a map $\mu:T_{k}\to S^{2}$ by
$\mu(e^{\eta i},ze^{\theta i})=\left\{\begin{array}[]{ll}(\cos(\eta/2),ze^{\theta i}\sin(\eta/2))_{R}&\ (0\leq\eta\leq\pi)\\ (-\cos(\eta/2),z\sin(\eta/2))_{R}&\ (\pi\leq\eta<2\pi).\end{array}\right.$
Then taking into account the fact that a principal circle bundle over $S^{1}$ is trivial we see that the classifying map of $p_{k}$ factors through $S^{2}$ and therefore we have
@@
$(Q_{k},p_{k},T_{k})\cong\mu^{*}(SU(2),p,S^{2})$ and $\mu^{*}:H^{2}(S^{2},\mathbb{Z})\stackrel{{\scriptstyle\cong}}{{\to}}H^{2}(T_{k},\mathbb{Z})$
$(1\leq k\leq m/2-1)$.
@@
In order to prove the first equation it suffices to show that there is a bundle map covering $\mu$. Let $\tilde{\mu}:Q_{k}\to SU(2)$ be the map given by
$\tilde{\mu}(D(e^{\eta i},ze^{\theta i})d(\bar{z}))=\left\{\begin{array}[]{ll}R(\bar{z}\cos(\eta/2),e^{\theta i}\sin(\eta/2))&\ (0\leq\eta\leq\pi)\\ R(-\bar{z}\cos(\eta/2),\sin(\eta/2))&\ (\pi\leq\eta<2\pi),\end{array}\right.$
which provides the required bundle map. The second equation is immediate from the definition of $\mu$. ∎
@@
$\psi(x,y)=p\biggl{(}\textstyle\prod_{j=0}^{m-2}R_{j}(r_{i;j},w_{i;j}))\textstyle\prod_{k=1}^{n-1}D(z_{2k-1},\zeta_{2k})\biggl{)}$
@@
forms the total space of a principal $S$-bundle endowed with the projection map $\tilde{q}:\tilde{P}\to(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ such that $\psi\circ\tilde{q}=p|P$. Let $\tilde{L}_{k}$ be the complex line bundle over $T_{k}$ associated to $p_{k}$. Write $\tilde{L}^{n-1}=\tilde{L}_{1}\boxtimes\cdots\boxtimes\tilde{L}_{n-1}$ and put $E=L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$. Then $(S(E),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong(\tilde{P},q,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})$ and therefore we have